# Can We Validate Counterfactual Estimations in the Presence of Unknown Network Interference?

### v1

@@ 6.1 LLM-based Social Network
This environment simulates a social media platform like Facebook and Quora where user interactions occur through content feeds, designed to study the effects of feed ranking algorithms on user engagement. The environment employs Large Language Model (LLM) agents to represent users, with demographically realistic personas derived from US Census data (U.S. Census Bureau, 2023) following Chang et al., (2024)’s methodology. Each agent possesses two interests selected from a predefined set of keywords: Technology, Sports, Politics, Entertainment, Science, Health, and Fashion.
@@
The treatment variable is the feed ranking algorithm, which determines content ordering for each user. In the control condition, users receive randomly ordered content, while the treatment condition presents content weighted by friend engagement. The outcomes of interest are user engagement metrics, measured as the sum of likes or replies generated by each user within a given time period, tracked in a comprehensive panel dataset.
@@
The underlying directed follower-following relationships network is constructed based on a preferential attachment model (Barabási and Albert, 1999). The dynamics of the environment center on content generation and user interactions. Content originates from an LLM-generated bank focused on a specified topic, with cross-keyword variations (e.g., climate change intersecting with technology or politics). The system generates feeds by combining interest-based content (matching user interests), trending content (high engagement), and random content according to specified proportions. Users interact with their feeds through an LLM-driven decision process that considers content relevance, friend engagement, feed position, and demographic characteristics.
@@
The simulation maintains comprehensive state information across multiple dimensions: engagement metrics, user interaction histories, network relationships, content visibility, and conversation threads. At each time step, agents have a 1% probability of generating new content, contributing to the platform’s organic content evolution. The interaction process follows a structured decision framework, where each agent evaluates content through a detailed prompt incorporating the agent’s persona attributes, content characteristics, social signals (including friend engagement), and feed positioning. This framework ensures that user behaviors and social influence patterns evolve naturally through the network while maintaining computational tractability.
6.2 Belief Adoption Model
@@
This environment models the diffusion of competing opinions within interconnected communities, implementing the cascade model of Montanari and Saberi, (2010). The system examines how opinions spread through social networks when individuals make decisions through coordination games with their neighbors. Through this framework, we evaluate the effectiveness of promotional campaigns in influencing opinion adoption patterns.
@@
The environment considers two competing stances: Opinion A (e.g., voting in an election) and Opinion B (e.g., declining to vote). The treatment represents a campaign aimed at increasing Opinion A adoption. The outcome for each unit in each period is binary: 1 if they adopt Opinion A, and 0 otherwise.
@@
The opinion evolution follows a network-based coordination game where each individual $i$ assigns payoff values $\text{payoff}^{i}_{\text{A}}$ and $\text{payoff}^{i}_{\text{B}}$ to both opinions. The probability of Opinion A adoption in the next period depends on the neighbor configuration and relative payoffs:
@@
$\mathbb{P}\big{(}\text{adopting Opinion~{}$\text{A}$}\big{|}n^{i}_{\text{A}},n^{i}_{\text{B}}\big{)}=\frac{e^{\beta\left(n^{i}h^{i}+n^{i}_{\text{A}}-n^{i}_{\text{B}}\right)}}{e^{\beta\left(n^{i}h^{i}+n^{i}_{\text{A}}-n^{i}_{\text{B}}\right)}+e^{-\beta\left(n^{i}h^{i}+n^{i}_{\text{A}}-n^{i}_{\text{B}}\right)}}\,,$
@@
where $h^{i}=\frac{\text{payoff}^{i}_{\text{A}}-\text{payoff}^{i}_{\text{B}}}{\text{payoff}^{i}_{\text{A}}+\text{payoff}^{i}_{\text{B}}}$ and $n^{i}_{\text{A}}$ represents the number of neighbors holding Opinion A out of $n^{i}$ total neighbors in the current period, and $\beta$ is a predetermined constant. The underlying network in our simulator is derived from the Pokec social network dataset (Takac and Zabovsky, 2012, Leskovec and Krevl, 2014), focusing on three regional networks: Krupina (3,366 users), Topolcany (18,246 users), and Zilina (42,971 users).
The environment also utilizes detailed profile data from the Pokec social network dataset to characterize each user. We extract three demographic variables from these profiles: age, profile activity, and gender. For base payoffs, we assign higher values for Opinion A to users aged 25-55 years and those maintaining active profiles. The treatment effectiveness follows a Gaussian distribution, reaching its maximum at age 35. Its impact scales directly with the user’s profile completion percentage, as measured in their Pokec data. These profile-driven calculations create distinct patterns of payoffs and treatment responses across the network, as visualized in Figures 2-4.
@@
Figure 2: Distribution of base payoffs, treatment effects, and node degrees for individuals in Krupina.
Figure 3: Distribution of base payoffs, treatment effects, and node degrees for individuals in Topolcany.
Figure 4: Distribution of base payoffs, treatment effects, and node degrees for individuals in Zilina.
6.3 Ascending Auction Model
This environment simulates a competitive market where multiple bidders participate in an ascending auction for objects, following the model of Bertsekas, (1990). The auction mechanism creates a dynamic pricing system where bidder interactions generate complex patterns of market influence, even without direct object-to-object relationships.
The system operates with $N$ objects and $N$ bidders. Each object represents an experimental unit, with its final value in each round serving as the outcome variable. The treatment consists of promotional interventions that increase bidder valuations by $\tau$for randomly selected objects. This treatment affects all bidders interested in the selected objects.
The market evolution follows a structured bidding process. In each round, bidders evaluate objects based on their private valuations and current market prices. They submit bids for their preferred objects, with objects being assigned to the highest bidders. These assignments establish new price levels, which influence subsequent bidding behavior. As prices increase through competitive bidding, objects become progressively less attractive to competing bidders.
The environment demonstrates a unique form of interference: while objects do not directly influence each other, treatment effects propagate through the market via bidders’ strategic responses to price changes. This creates a network of indirect treatment effects, as promotional interventions for certain objects can influence market outcomes for others through shifts in bidder behavior and price dynamics.
6.4 New York City Taxi Routes
@@
This environment models ride-sharing dynamics across New York City taxi zones using real-world TLC Trip Record Data (New York City Taxi and Limousine Commission, 2024). The framework adapts the established linear-in-mean outcome model (Eckles et al., 2016, Cai et al., 2015, Leung, 2022) to represent how passengers utilize ride-sharing services throughout the city. By incorporating actual travel data, passenger density metrics, and inter-zone relationships, the simulation effectively captures the complex network of interactions between taxi routes across the city.
@@
In this setting, the experimental units are defined as routes (origin-destination pairs) between the city’s 263 taxi zones, with time segmented into 6-hour periods. The outcome variable measures the number of trips along each route during each time period. The treatment represents a program implemented on randomly selected routes and the goal is to evaluate travelers response patterns.
@@
Given baseline outcomes $[y^{i}_{t}{}]_{i,t}$, the system’s evolution follows Equation (12), where outcomes depend on baseline patterns, network effects, and treatment status:
@@
Figure 5: Mean and variance of the number of trips in each route during each period, revealing a strong daily and weekly seasonality pattern.
The environment incorporates real-world data through three components: First, it uses “High Volume For-Hire Vehicle Trip Records” (January-March 2024, 57,974,677 trips) to construct baseline outcomes (denoted by $[y^{i}_{t}{}]_{i,t}$), focusing on 18,768 active routes across 366 periods. As shown in Figure 5, the trips have a strong seasonality pattern, which is common in the ride-hailing application (Xiong et al., 2024b ). Second, it employs the “Taxi Zone Lookup Table” and Claude (Model 3.5 Sonnet, Anthropic, 2024) to generate passenger density scores that determine route-specific treatment effects (Figure 6 left). Third, it creates a route adjacency network based on geographic proximity, transit connections, shared roads, and functional relationships, yielding an average node degree of 8.32 (Figure 6 right). This data-driven approach ensures the simulation reproduces key real-world characteristics: temporal patterns, heterogeneous treatment effects, and localized network interactions.
@@
Figure 6: The left panel displays the distribution of route-specific direct treatment effects, illustrating the heterogeneity in treatment responses. The right panel shows the histogram of node degrees in the route interference network.
6.5 Exercise Encouragement Program
This environment simulates an exercise intervention program that combines individual characteristics from the 1994 Census Bureau database (Kohavi and Becker, 1994) with social network effects. Drawing inspiration from mobile health intervention studies (Liao et al., 2016, Klasnja et al., 2015, 2019), the environment models how digital encouragement messages influence exercise decisions within a social network context.
@@ 6.5 Exercise Encouragement Program
The experimental units are individuals, with binary outcomes representing their exercise decisions in each period (1 for exercise, 0 for no exercise). The treatment consists of digital intervention messages designed to encourage physical activity. Inspired by Li and Wager, 2022a , we let the outcomes follow a Bernoulli distribution defined as follows:
@@
$Y^{i}_{t+1}{}\sim\text{Bernoulli}\left(\frac{1}{1+\exp{-(\delta_{t}^{i}+\tau_{t}^{i}W^{i}_{t+1}+\theta Y^{i}_{t}{}Z^{i}_{t}+\lambda W^{i}_{t+1}Y^{i}_{t}{}Z^{i}_{t})}}\right),$ (13)
@@
where $Z^{i}_{t}=\sum_{j=1}^{N}E^{ij}Y^{j}_{t}{}$ represents the count of neighboring individuals who exercised in the previous period. Here, $E^{ij}$’s are the elements of the adjacency matrix from Twitter social circles data (Leskovec and Mcauley, 2012), with an average of 21.74 connections per individual (Figure 9).
In (13), each component captures specific aspects of exercise behavior. The baseline probability ($\delta_{t}^{i}$) represents an individual’s inherent tendency to exercise, derived from Census Bureau demographic data. This probability incorporates age (with higher values for younger individuals), working hours (showing an inverse relationship with exercise likelihood), and occupation type (assigning higher probabilities to active or professional occupations). These baseline probabilities exhibit weekly patterns, showing peak values on weekends due to increased free time, elevated rates on Mondays from new week motivation, stable mid-week patterns, and slightly lower values on Fridays reflecting end-of-week fatigue (Figure 7).
The intervention effectiveness ($\tau_{t}^{i}$) quantifies how individuals respond to exercise encouragement messages. This response varies based on multiple demographic factors from the Census Bureau database. Younger individuals show higher responsiveness to digital interventions, while education level correlates positively with intervention effectiveness. Job-related factors, including occupation type and working hours, influence response rates by indicating flexibility and availability to act on interventions. The impact of messages follows weekly cycles, demonstrating maximum effectiveness during weekends and Mondays, with gradually decreasing impact through mid-week (Figure 8). The model also sets parameters $(\theta,\lambda)=(0.0.4,0.01)$ to capture peer influence and the interaction between treatment and peer effects.
Figure 7: Distribution of baseline exercise probabilities across the population.
Figure 8: Distribution of intervention message effects across the population
@@
6.6 Data Center Server Utilization
@@
This environment simulates a server farm, and the goal is to evaluate interventions for improving server utilization. Given the increasing demand for cloud computing resources, optimizing data center utilization has become critical for addressing global sustainability concerns (Zhang et al., 2023, Saxena et al., 2023). Within this system, servers influence each other’s performance through the system’s physical characteristics, particularly via the join-the-shortest-queue routing policy (Gupta et al., 2007). This operational dynamic creates an implicit interference pattern in the absence of a pre-specified network structure.
The experimental units are individual servers within a parallel processing system of N servers. The outcome variable $Y^{i}_{t}{}$ represents server i’s utilization during the interval [t,t+1), measured as the proportion of time the server remains busy. The treatment consists of interventions that enhance the processing power of selected servers.
@@
The system evolution follows a structured routing mechanism. When tasks arrive, the system identifies capable servers for each task type and selects a random sample among them. Following the join-the-shortest-queue policy, tasks are assigned to servers with minimal queue lengths within this sample, using random assignment to resolve ties. This routing approach naturally creates interference effects, as performance improvements in treated servers influence task distribution across the entire system.
@@
The environment incorporates realistic workload patterns through a time-dependent Poisson arrival process. The demand model captures multiple temporal patterns: daily variations (night-time lows, morning increases, midday peaks, and evening declines), weekly cycles (heightened weekday activity), and stochastic elements (random fluctuations and event-driven spikes). Each server processes tasks with exponentially distributed service times, which can be modified by interventions. Through this design, the system replicates key characteristics of real-world data centers while enabling controlled experimentation.
@@
7 Results
Figures 11-18 present results from applying counterfactual cross-validation (Algorithm 2) across six benchmark scenarios detailed in §6. Below, we outline our implementation approach and key findings.

### v2

@@ 6.1 LLM-based Social Network [Uncaptioned image]
This environment simulates a social media platform like Facebook or Quora where user interactions occur through content feeds, designed to study the effects of feed ranking algorithms on user engagement. The environment employs Large Language Model (LLM) agents to represent users, with demographically realistic personas derived from US Census data (U.S. Census Bureau, 2023) following Chang et al., (2024)’s methodology. Each agent possesses two interests selected from a set of keywords: Technology, Sports, Politics, Entertainment, Science, Health, and Fashion.
@@
The treatment variable is the feed ranking algorithm, which determines the ordering of content presented to each user. In the control condition, users receive content in random order, while in the treatment condition, content is ranked by weighting based on friend engagement levels. The outcome of interest is user engagement metrics, measured as the total number of likes and replies by each user within a time period. The network structure is then captured by a directed follower-following graph constructed using a preferential attachment model (Barabási and Albert, 1999).
@@
The dynamics of the environment center on content generation and user interactions. Content originates from a bank focused on a specified topic (e.g., climate change) with cross-keyword variations (e.g., technology or politics). The system then generates user-specific feeds by combining interest-based content (matching user interests), trending content (high engagement), and random content. In each round, users interact with their feeds through an LLM-driven decision process that incorporates content relevance, friend engagement, feed position, and demographic characteristics.
@@
The simulation maintains comprehensive state information including engagement metrics, user interaction histories, network relationships, content visibility, and conversation threads. At each time step, agents have a 1% probability of generating new content, contributing to organic content evolution. This interaction framework ensures that user behaviors and social influence patterns evolve naturally through the network while maintaining computational traceability.
Figure 6: LLM-based Social Network. Demographically diverse AI agents interact through content feeds. Treatment involves optimized vs. random feed ranking, with outcomes measured as user engagement metrics.
6.2 Belief Adoption Model [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
This environment models the diffusion of competing opinions within interconnected communities, implementing the cascade model of Montanari and Saberi, (2010). The setting examines how opinions spread through social networks when individuals make decisions through coordination games with their neighbors. Through this framework, we evaluate the effectiveness of promotional campaigns in influencing opinion adoption patterns.
@@
The environment considers two competing stances: Opinion A (e.g., voting in the upcoming election) and Opinion B (e.g., not voting in the election). The treatment represents a campaign aimed at increasing Opinion A adoption. The outcome for each unit in each period is binary: 1 if they adopt Opinion A, and 0 otherwise.
@@
The opinion evolution follows a network-based coordination game where each individual $i$ assigns base payoff values $\text{payoff}^{i}_{\text{A}}$ and $\text{payoff}^{i}_{\text{B}}$ to both opinions. The probability of Opinion A adoption in the next period depends on the neighbor configuration and relative payoffs:
@@
$\mathbb{P}\big(\text{adopting Opinion~$\text{A}$}\big|n^{i}_{\text{A}},n^{i}_{\text{B}}\big)=\frac{1}{1+e^{-2\beta\left(n^{i}h^{i}+n^{i}_{\text{A}}-n^{i}_{\text{B}}\right)}},\quad\quad h^{i}=\frac{\text{payoff}^{i}_{\text{A}}-\text{payoff}^{i}_{\text{B}}}{\text{payoff}^{i}_{\text{A}}+\text{payoff}^{i}_{\text{B}}},$
@@
where $n^{i}_{\text{A}}$ represents the number of neighbors holding Opinion A out of $n^{i}$ total neighbors in the current period, and $\beta$ is a predetermined constant. The underlying network in our simulator is derived from the Pokec social network dataset (Takac and Zabovsky, 2012, Leskovec and Krevl, 2014), focusing on three regional networks: Krupina (3,366 users), Topolcany (18,246 users), and Zilina (42,971 users). For treated individuals, the final payoff equals the base payoff plus the treatment effect, where both components are characterized for each user through three key demographic variables: age, profile activity, and gender from the Pokec social network dataset.
Individual Level Demographics: Age and profile activity $\downarrow$ Base payoffs and treatment effect $\downarrow$ Heterogeneous responses Network Level Network effects amplify individual changes through coordination games with the neighbors $\downarrow$ Opinion cascades through social connections
Figure 7: Belief Adoption Model. Individual demographics determine heterogeneous base payoffs and treatment responses, which then propagate through network to create opinion cascades across the network.
Base payoffs for Opinion A are strategically calibrated to reflect realistic demographic patterns. Users aged 25-55 years receive elevated base payoffs for Opinion A, capturing the tendency for individuals in this age range to exhibit greater civic engagement. Furthermore, users with higher profile activity levels, quantified by the completeness of their profile information, are assigned increased base payoffs for voting. This reflects the fact that socially engaged users typically demonstrate heightened civic participation and political awareness. These demographic adjustments create heterogeneous baseline voting propensities across the network, with distinct population segments exhibiting varying predispositions toward electoral participation.
Treatment effectiveness employs a complementary modeling approach. The treatment effect follows a Gaussian distribution centered at age 35, representing peak responsiveness to new content (e.g., get-out-the-vote messaging) within the adult population. Complementing these age-based effects, treatment impact also scales proportionally with each user’s profile activity level. Users maintaining more comprehensive profiles (indicative of greater platform engagement and social connectivity) experience amplified treatment effects. This captures the phenomenon that active social media participants exhibit increased susceptibility to civic campaigns and peer-mediated influence. This approach generates realistic heterogeneity in treatment responsiveness across the network, as demonstrated in Figure 8.
@@
Figure 8: Distribution of base payoffs, treatment effects, and node degrees for individuals across three regions: Krupina with 3,366 users (top), Topolcany with 18,246 users (middle), and Zilina with 42,971 users (bottom).
6.3 New York City Taxi Routes [Uncaptioned image]
@@
This environment is based on ride-sharing dynamics across New York City taxi zones using TLC Trip Record Data (New York City Taxi and Limousine Commission, 2024). The framework adopts the established linear-in-means outcome specification (Eckles et al., 2016, Leung, 2022) to model passenger response patterns to a pricing experimentation, while incorporating realistic noise levels, seasonality patterns, and temporal correlations observed in the actual data. By leveraging actual trip records, passenger density metrics, and inter-zone connectivity, the simulation captures the complex network of interactions between taxi routes throughout the city across different periods.
Temporal Variations Real seasonality: • Daily rush patterns • Weekly demand cycles Data scale: 58M taxi trips in NYC Complex Interference Route connections: • Geographic proximity • Shared infrastructure Network structure: 18,768 routes, avg degree: 8.32 Response Heterogeneity Route differences: • Passenger density • Location characteristics Treatment effects: AI-generated route-specific scores
Figure 9: NYC Taxi Model. The environment incorporates authentic temporal patterns from 58M trips, realistic network interference through route adjacency, and heterogeneous treatment responses.
@@
In this setting, we conceptualize the treatment as a new pricing algorithm that incorporates route-specific characteristics (e.g., distance, average demand and supply at origin and destination). Each route corresponds to an origin-destination pair between the city’s 263 taxi zones. The new pricing algorithm is then implemented on randomly selected routes to evaluate travelers’ response patterns. We segment time into 6-hour periods, with the outcome variable measuring the number of trips along each route during each period.
@@
Precisely, given baseline outcomes $y^{i}_{t}{}]_{i,t}$, the system’s evolution follows Equation (17):
@@
Figure 10: Mean and variance of baseline outcomes $y_{t}^{i}$ (the number of trips on route $i$ during period $t$), revealing strong daily and weekly seasonality patterns.
@@
Figure 11: Distribution of route-specific direct treatment effects (left) and node degree distribution in the route interference network (right), both illustrating heterogeneity in treatment responses and local interactions.
In (17), the outcome $Y^{i}_{t}{}$ represents the number of trips over route $i$ during period $t$ under a specific treatment allocation. To naturally reflect potential real-world outcome variations both across units and over time, we extract baseline outcomes $y^{i}_{t}{}$ from “High Volume For-Hire Vehicle Trip Records” for January-March 2024 (57,974,677 trips), resulting in 18,768 active routes across 366 periods. As shown in Figure 10, the baseline outcomes demonstrate strong seasonality patterns, which are common in ride-hailing applications (Xiong et al., 2024b ).
The adjacency network captured by $\bf E$} in (17) is constructed using AI models (Claude 3.5 Sonnet, Anthropic, 2024) based on the “Taxi Zone Lookup Table.” Specifically, the adjacency of two routes is determined according to geographic proximity, transit connections, shared roads, and functional relationships, yielding a network with an average node degree of 8.32 (Figure 11, right panel). Finally, route-specific direct treatment effects $\tau_{\text{u}}^{i}$ are generated using passenger density scores in both origin and destination, also determined by AI models (Figure 11, left panel). This approach is to ensure the simulation reproduces key real-world temporal variations and noise patterns.
6.4 Exercise Encouragement Program [Uncaptioned image]
This environment models a digital health intervention program designed to assess how sending digital encouragement messages influences users’ exercise decisions (Liao et al., 2016, Klasnja et al., 2015, 2019). To enhance realism, we integrate individual-level features from the Census Bureau database (Kohavi and Becker, 1994) with Twitter social circles data (Leskovec and Mcauley, 2012). This combination allows us to capture user diversity and their heterogeneous social interactions.
Figure 12: Exercise Encouragement Program. Individual decisions depend on personal characteristics, treatment status, and friends’ exercise patterns, creating dynamic loops where exercise behavior spreads through social ties.
@@ 6.4 Exercise Encouragement Program [Uncaptioned image]
In this setting, the experimental units are individuals, with binary outcomes representing their exercise decisions in each period (1 for exercise, 0 for no exercise). The treatment consists of sending digital intervention messages designed to encourage physical activity. Inspired by Li and Wager, 2022a , we let the outcomes follow a Bernoulli distribution defined as follows:
@@
$Y^{i}_{t+1}{}\sim\text{Bernoulli}\left(\frac{1}{1+\exp{[-(\delta_{t}^{i}+\tau_{t}^{i}W^{i}_{t+1}+\theta Y^{i}_{t}{}Z^{i}_{t}+\lambda W^{i}_{t+1}Y^{i}_{t}{}Z^{i}_{t}-\eta V_{i}^{t})}]}\right),$ (18)
@@
where $Z^{i}_{t}=\sum_{j=1}^{N}E^{ij}Y^{j}_{t}{}$ represents the count of neighboring individuals who exercised in the previous period, and $V_{t}^{i}$ represents the variance of exercise behavior among individual $i$’s neighbors in the previous period. Here, $E^{ij}$’s are the elements of the adjacency matrix from Twitter social circles data, with an average of 21.74 connections per individual (Figure 13).
@@
In Equation (18), each component captures distinct aspects of exercise behavior. The baseline score ($\delta_{t}^{i}$) represents an individual’s inherent propensity to exercise, uniquely determined for each participant based on Census Bureau demographic data. This score reflects factors such as age (with younger individuals assigned higher values), working hours (which show an inverse relationship with exercise likelihood), and occupation type (favoring active or professional occupations). These baseline scores exhibit pronounced weekly patterns: they peak on weekends due to increased free time, rise on Mondays from renewed weekly motivation, stabilize mid-week, and slightly decline on Fridays, reflecting end-of-week fatigue (Figure 14).
Figure 14: Distribution of baseline exercise score $\delta_{t}^{i}$ across the population over weekdays.
The intervention effectiveness ($\tau_{t}^{i}$) quantifies how individuals respond to digital messages. This response varies based on multiple demographic factors. Younger individuals show higher responsiveness to digital interventions and education level correlates positively with intervention effectiveness. Job-related factors, such as occupation type and working hours, further influence response rates by reflecting flexibility and availability to act on the interventions. Moreover, the impact of messages follows a weekly cycle: effectiveness peaks on weekends and Mondays, then gradually declines through mid-week (Figure 15). The model also incorporates parameters $(\theta,\lambda,\eta)=(0.04,0.01,0.02)$ to capture various types of peer influence effects.
Figure 15: Distribution of intervention message effects across the population over weekdays.
6.5 Data Center Server Utilization [Uncaptioned image]
@@
This environment simulates a server farm to evaluate interventions designed to improve server utilization. With the increasing demand for cloud computing resources, optimizing data center utilization has become critical for addressing global sustainability concerns (Zhang et al., 2023, Saxena et al., 2023). Within this system, servers influence each other’s performance through the system’s physical characteristics, particularly via the task routing policy. This operational dynamic creates an implicit interference pattern in the absence of a pre-specified explicit network structure.
Figure 16: A simple four-server data center. Join-the-shortest-queue routing creates implicit interference: Server 2 with enhanced processing power (treatment) maintains consistently shorter queues, causing the routing system to direct more tasks to this server, reducing available demand for untreated servers.
@@
The experimental units in this setting are individual servers in a parallel processing system of $N$ servers. The outcome variable $Y^{i}_{t}{}$ represents server $i$’s utilization during the interval $[t,t+1)$, measured as the proportion of time the server is busy. The treatment consists of interventions that enhance the processing power of selected servers. Then, the system evolves according to a structured routing mechanism. When tasks arrive, the system identifies capable servers for each task type and randomly samples among them. Following the join-the-shortest-queue policy (Gupta et al., 2007), tasks are assigned to the servers with the shortest queues within this sample, with random assignment used to break ties. This routing policy naturally generates interference effects, as performance improvements in treated servers shift task distribution across the system (Figure 16).
@@
The environment captures realistic workload patterns using a time-dependent Poisson arrival process. This demand model features daily cycles (nighttime lows, morning ramp-ups, midday peaks, and evening declines) overlaid with stochastic fluctuations to mimic real-world variability (Figure 17). Beyond these dynamic patterns, tasks are categorized into multiple types, each with its own compatibility constraints. Servers are also assigned a subset of job types they can process, creating a diverse landscape of server capabilities. Each server processes tasks with exponentially distributed service times, where the processing rates can be enhanced by treatment interventions. Collectively, these mechanisms replicate critical aspects of real-world data centers, such as dynamic routing and resource contention, while enabling controlled experimentation and measurement of treatment effects.
@@
6.6 Auction Model [Uncaptioned image]
This environment simulates a stylized auction model with explicit interference through a bipartite structure. In this setting, multiple bidders participate in an auction for objects, following the model of Bertsekas, (1990). The auction mechanism creates a simplified pricing system where bidder interactions generate basic patterns of market influence without direct object-to-object relationships.
The setting operates with $N$ objects and $N$ bidders. Each object represents an experimental unit, with its final value in each period serving as the outcome variable. The treatment consists of promotional interventions that increase bidder valuations by $\tau$% for randomly selected objects. This treatment affects all bidders interested in the selected objects.
The market evolution follows a simple bidding process. In each round, bidders evaluate objects based on their private valuations and current market prices. They submit bids for their preferred objects, with objects being assigned to the highest bidders. These assignments establish new price levels, which influence subsequent bidding behavior. As prices increase through competitive bidding, objects become progressively less attractive to competing bidders. This leads to a unique form of interference: while objects do not directly influence each other, treatment effects propagate through the market via bidders’ strategic responses to price changes. This creates a network of indirect treatment effects, as promotional interventions for certain objects can influence market outcomes for others through shifts in bidder behavior and price dynamics (Figure 18).
To enhance realism, the environment implements different bidder types with distinct valuation patterns: standard bidders provide baseline valuations, collectors value items higher with more variability, dealers maintain slightly lower but more consistent valuations, and investors exhibit higher but more variable valuations. This approach captures how different market participants value the same items differently, creating heterogeneous pricing strategies.
Figure 18: Auction model for two objects in two bidding rounds. Upper panel: base scenario with no treatment. Lower panel: treating Object A with a 10% valuation increase raises its final value from \$100 to \$130.9, while Object B’s value decreases from \$90 to \$70, demonstrating indirect market interference effects.
7 Results and Discussions
Table 2 provides implementation details for our experimental setup across six diverse scenarios described in §6, with results presented in Figures 1925. Each experimental run executes one iteration of the respective setting, followed by applying our counterfactual cross-validation framework (Algorithm 2) for estimation and validation. We outline our implementation strategy below and summarize the key empirical findings.
Common Implementation Parameters
Initial Condition All scenarios start from $t=0$ with no treatment
Treatment Design Staggered rollout over three stages
Baseline Methods basic Causal Message-Passing (bCMP)
Difference-in-Means (DM), Horvitz-Thompson (HT)
Validation Setup for Counterfactual Cross-Validation (Algorithm 2)
Validation Batches $b_{v}=2$ batches
Time Blocks Three equally-length blocks
Number of Batches ($\mathsf{n_{b}}$) 100, 500, 1000
Batch Sizes ($\mathsf{s_{b}}$) 5%, 10%, 20%, 30%, 50% of population NN
Model Ridge regression with $\alpha\in\{10^{-4},10^{-2},1,100\}$
Scenario-Specific Implementation Details
Scenario Population ($N$) Each Stage Length Runs Design $\vec{p}$
LLM Social Network 1,000 10 (total: $T=31$) 10 $(0.2,0.5,0.8)$
Belief Adoption - Krupina 3,366 2 (total: $T=7$) 100 $(0.1,0.2,0.5)$
Belief Adoption - Topolcany 18,246
Belief Adoption - Zilina 42,971
NYC Taxi Routes 18,768 28 (total: $T=85$)
Exercise Encouragement 30,162 7 (total: $T=22$)
Data Center 2,000 24 (total: $T=73$)
Auction Model 200 5 (total: $T=16$)
Table 2: Implementation Details